determine both the mass and the charge of a block of material consisting of 3 x 10 ^27 protons,3 x10^37 neutrons, and 3.1 x 10^27 electrons


1
Expert's answer
2022-05-10T13:21:19-0400

The mass:

"m=N_pm_m+N_nm_n+N_em_e\\
=3*10^{27}*1.67*10^{-27}+3*10^{37}*1.67*10^{-27}\\
+3.1*10^{27}*9.1*10^{-31}=5.01*10^{10}\:\rm kg"

The caharge:

"q=N_pq_m+N_eq_e\\
=3*10^{27}*1.6*10^{-19}
-3.1*10^{27}*1.6*10^{-19}\\
=-1.6*10^{7}\:\rm C"